A bank loaned out php12.000, part of it at the rate of 8%per yaer and the rest at the rate of 18%per year. If the invest recieved totaled php1,000, how much was loaned at 8%
Answer by stanbon(75887)

A bank loaned out php12.000, part of it at the rate of 8%per yaer and the rest at the rate of 18%per year. If the invest recieved totaled php1,000, how much was loaned at 8%
---------------
Let amount invested at 8% be "x" ; interest on that is 0.08x dollars
---------
Amt invested at 18^ is "12,000-x"; int is 0.18(1200-x)= 2160-0.18x dollars
------------------
EQUATION:
interest + interest = 1000
0.08x + 2160-0.18x = 1000
-0.10x = -1160
x = 11,600 (amount invested at 8%)
12,000-11600 = 400 (amount invested at 18%)
===============
Cheers,
Stan H.